A super confusing word problem to an average math student.
"A laboratory used 120 fence posts in an experiment comparing two types of paint. Six fewer than twice as many fence posts were painted with paint A as were painted with paint B. How many fence posts were painted with paint A? Paint B?"
Would be glad to get your help!

Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

a=2b-6
a+b=120
2b-6+b=120
3b=126
b=42
a=78
Paint A was used on 78 fence posts; paint B was used on 42 posts..

Answer by stanbon(75887)

A laboratory used 120 fence posts in an experiment comparing two types of paint.
------------
Six fewer than twice as many fence posts were painted with paint A as were painted with paint B.
-----------------
How many fence posts were painted with paint A? Paint B?"
====
Equations:
A + B = 120 posts
A = 2B -6
------------
Subtract for "A" and solve for "B":
2B-6 + B = 120
3B = 126
B = 42 (# of B-painted posts)
----
Solve for "A":
A = 2B-6
A = 2*42-6
A = 84-6
A = 76 (# of A-painted posts)
==============================
Cheers,
Stan H.
==============